OpenStudy (donnie1999):

HELLP MEEE PLEASEE!!!!! find the slope and the y-intercept of the equation y-3(x-1)=0

OpenStudy (johnweldon1993):

The slope-intercept form of a line is \[\large y =mx + b\] So we need to make our expression look like that \[\large y - 3(x - 1) = 0\] Where would be start?

OpenStudy (donnie1999):

find what x is equal to I guess

OpenStudy (johnweldon1993):

Not nearly as complicated, all we need to do is rearrange our equation so that y = something since y = mx + b is the form we want So from \[\large y-3(x-1) = 0\] the first step would be to distribute that 3 into the parenthesis...we want to simplify this a bit

OpenStudy (donnie1999):

so it would be 3x-3

OpenStudy (johnweldon1993):

Well yes it could be...normally we would distribute the minus sign with it but that's okay..but we have to make sure that now we have \[\large y - (3x - 3) = 0\] right? What would be next...remember we want y = ...

OpenStudy (donnie1999):

so do I multiply 0*3x or 0*3

OpenStudy (johnweldon1993):

No we dont need to multiply at all here \[\large y - (3x - 3) = 0\] If the only thing stopping us from having that y = something is a \(\large -(3x - 3)\) then all we have to do is add it to the other side \[\large y \cancel{- (3x - 3) + (3x - 3)} = 0 + (3x - 3)\] \[\large y = ?\]

OpenStudy (donnie1999):

so y=-3 srry im kind of clueless

OpenStudy (johnweldon1993):

No dont worry about it algebra is always tricky in the beginning, actually depending on what you just asked then you're right or wrong lol Well....if we have \(\large y = 0 + (3x - 3)\) that just means we have \[\large y = 3x - 3\] right? So the slope would be? the y-intercept would be?

OpenStudy (donnie1999):

ok so the slope would be 3 and the y-intercept is -3

OpenStudy (johnweldon1993):

Perfect!